5) A company has systematic risk (or beta factor) 1.25, the risk-free rate of interest is currently 5.45%, and the required return on the market portfolio is 12%. The company plans to pay a dividend of AED 3.05 per share in the coming year (2020) and anticipates that its future dividends will increase at an annual rate consistent with that experienced over the 2001-2003 period:

End of Year     Dividend

2017                     2.00

2018                     3.00

2019                     2.10

Estimate the stock price value of the company’s share using CAPM.

Q2. (i). Explain the management the risks and opportunities given contemporary business conditions (Hint: upward slope of the yield curve). What is your advice to the firm? (ii). How demand and supply bond curves react under these conditions?

Q3. (i). How is the value of a bond determined? (ii). Which factors affect price sensitivity?

Q4. What is the value of a 5-year, AED 1,000 par value bond with an 8% annual coupon if the YTM (required rate of return) of return is 8%? Explain to the manager of the firm without calculation.

If you use excel do show formulas
Answer #1

You have asked multiple unrelated questions in the same post. I have addressed the first one. Please post the balance questions separately one by one.

Q - 5

CAGR in dividend, g = (D2019 / D2017)1/2 - 1 = (2.10 / 2.00)1/2 - 1 = 2.47%

D2020 = 3.05

Cost of equity = Ke = Rf + Beta x (Rm - Rf) = 5.45% + 1.25 x (12% - 5.45%) = 13.64%

Hence, the stock price value of the company’s share using CAPM = D2020 / (Ke - g) = 3.05 / (13.64% - 2.47%) = AED 27.31